OpenStudy (anonymous):

**Please help!! Which of the following is irrational? square root of ten square root of twenty-five square root of one hundred square root of two hundred twenty-five

hartnn (hartnn):

simple questions : \(5 \times 5 = ... \\ 10 \times 10 = ... \\ 15 \times 15 = ...\)